OpenStudy (anonymous):

ANSWER IN 5 MINUTES FOR MEDAL + FAN!! Simplify. - |-5.2|

OpenStudy (anonymous):

-5.2 is my answer

OpenStudy (igreen):

The absolute value will change it to -(5.2). Distributing the -1, will change it back to -5.2.

OpenStudy (anonymous):

-5.2 it is

OpenStudy (anonymous):

mine to -5.2

OpenStudy (anonymous):

cool i got it first, doesnt happen very often :P

OpenStudy (anonymous):

-5.2

OpenStudy (tkhunny):

Why do we get only five minutes? Are we cheating on a timed exam?

OpenStudy (morganchavez):

no its jus lik bc we have more work to do and dont have allday for somone to finally answer 3 hours later

OpenStudy (anonymous):

XD yep thats how i feel

OpenStudy (morganchavez):

bc that has happened to me it took me a week to try n get some real help

OpenStudy (morganchavez):

and if we dont put tht then we never get help at all :)

OpenStudy (anonymous):

i need help please Unit 3 | Lesson 6 Lesson Checkpoint: Show What You Know Question Navigator 4. Which equation can be used to find the answer? Joelle's stamp book has room for 54 stamps. She already has 38 stamps in the book. How many more stamps does she need to fill the book? A. s – 38 = 54 s = 92 92 stamps B. 2 · s = 54 s = 27 27 stamps C. s + 38 = 54 s = 16 16 stamps D. s ÷ 2 = 38 s = 76 76 stamps

OpenStudy (anonymous):

well look at all the ppl she got XD

OpenStudy (morganchavez):

c point proven true

OpenStudy (tkhunny):

@morganchavez I see, so your time is more valuable than the time of the volunteers who so nicely do the best they can to answer your questions. Seriously, if you wish instant responses and snappy results, you should consider paying for them.

OpenStudy (anonymous):

minus for a times for c and division for d

OpenStudy (anonymous):

Hold on @J-bird. I'll fan u it's just that my network is a ll wack

OpenStudy (anonymous):

mine is doing that too dont know y but its hard to fan

OpenStudy (tkhunny):

@Devilwolf3000 Why are we hijacking threads? Start your own and SHOW YOUR WORK.

OpenStudy (anonymous):

mine too

OpenStudy (anonymous):

wat?

OpenStudy (igreen):

@tkhunny Everyone on this site who asks a question cheats.

OpenStudy (anonymous):

hijacking threads what is that

OpenStudy (igreen):

There's no changing that. Where do you think they get their questions from? Obviously taken from their tests.

OpenStudy (anonymous):

true true

OpenStudy (morganchavez):

ummm excuse me i didnt mean d=to offend u but everyone always skipps pple and they post things on ur posts lik oooo hi n ur like hello n then they say so how old r u lik excuse me this aint no dating site ok i think i kow what i am saying @tkhunny

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

and @tkhunny he is right it is TOOOOO MUUUUCHHHHHH WORKKKK TOOOO DOOOOOOO WEEEEE GETTTTT TIREEEDDDDDDD AFTERRRRRR A WHOLEEEEEEE DAYYYYY OFFFFFFF WORRRRKKKKKKK MS.I WANNA TATTLETALE

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

yeah stay of our backs @tkhunny

OpenStudy (anonymous):

what does XD means

OpenStudy (anonymous):

I WILL fan u @J-bird I SWEAR I'm just having a hard time doin that

OpenStudy (anonymous):

I feel really bad but I will eventually

OpenStudy (anonymous):

*coughs at tkhuuny syain NUN*

OpenStudy (anonymous):

Which equation can be used to find the answer? Clarice spent $27 of her savings on clothes and has $49 of savings left. How much did she have in savings before buying the clothes? A. m + 27 = 49 m = 22 $22 B. m – 27 = 49 m = 76 $76 C. m ÷ 2 = 27 m = 54 $54 D. m ÷ 2 = 49 m = 98 $98

OpenStudy (anonymous):

everyone is having this problem today dont worry about it

OpenStudy (anonymous):

yeah and what does that stupid nun have against us right @everyone

OpenStudy (igreen):

@tkhunny is a Moderator..she doesn't have to tattletale..she can just ban you.

OpenStudy (igreen):

He/She..

OpenStudy (anonymous):

i dont care

OpenStudy (anonymous):

he or she can suck it if all i care

OpenStudy (tkhunny):

@iGreen It is false that everyone cheats. Perhaps there are many. Perhaps everyone you know. Nonetheless, your statement is false. Even if it were true, YOU still would have the option to maintain your integrity and NOT cheat. It is a choice. No one doing anything excuses your behavior or forces you to do anything. You can choose to be honest.